# MedPT: A Massive Medical Question Answering Dataset for Brazilian-Portuguese Speakers

### v1

@@ Abstract
While large language models (LLMs) show transformative potential in healthcare, their development remains focused on high-resource languages, creating a critical barrier for others as simple translation fails to capture unique clinical and cultural nuances, such as endemic diseases. To address this, we introduce MedPT, the first large-scale, real-world corpus for Brazilian Portuguese, comprising 384,095 authentic question-answer pairs from patient-doctor interactions. The dataset underwent a meticulous multi-stage curation protocol, using a hybrid quantitative-qualitative analysis to filter noise and contextually enrich thousands of ambiguous queries. We further augmented the corpus via LLM-driven annotation, classifying questions into seven semantic types to capture user intent. Our analysis reveals its thematic breadth (3,200 topics) and unique linguistic properties, like the natural asymmetry in patient-doctor communication. To validate its utility, we benchmark a medical specialty routing task: fine-tuning a 1.7B parameter model achieves an outstanding 94% F1-score on a 20-class setup. Furthermore, our qualitative error analysis shows misclassifications are not random but reflect genuine clinical ambiguities (e.g., between comorbid conditions), proving the dataset’s deep semantic richness. We publicly release MedPT to foster the development of more equitable, accurate, and culturally-aware medical technologies for the Portuguese-speaking world.
@@
MedPT: A Massive Medical Question Answering Dataset for Brazilian-Portuguese Speakers
@@ 1. Introduction
Large language models (LLMs) are increasingly demonstrating transformative potential across diverse healthcare applications, from online patient support and clinical text classification to robust medical question answering Nazi and Peng (2024); Liu et al. (2025). However, the performance, safety, and clinical reliability of these systems are fundamentally contingent on the scale and quality of their training data Ryan Conmy et al. (2023). Consequently, datasets that lack demographic representation, contain inherent biases, or are insufficiently curated to handle complex medical nuances can lead to models that produce factually incorrect or systematically skewed outputs, posing significant risks and undermining clinical trust Khan et al. (2021); Schwabe et al. (2024). Compounding this challenge, current progress and resource development are overwhelmingly concentrated in high-resource languages, particularly English, a linguistic bias that creates a critical barrier for other major world languages, severely limiting the development, validation, and equitable deployment of these promising technologies to large global populations Enshaei and Naderkhani (2024).
This linguistic and data-centric gap is clearly reflected in the literature, where significant research investment in medical text curation has centered on English Jin et al. (2019); Pal et al. (2022); Singhal et al. (2022); Preston et al. (2023), Chinese He et al. (2019); Zhang et al. (2018); Zhu et al. (2023), or both He et al. (2020). Despite valuable initiatives have begun to extend this scope to other languages such as Arabic, Vietnamese, Persian, and Turkish Abdelhay et al. (2023); Nguyen et al. (2022); Ghassabi et al. (2025); İncidelen and Aydoğan (2024), the resource disparity remains particularly acute for Brazilian Portuguese in specialized domains such as open-domain medical question answering.
@@
To address this critical resource gap, we introduce MedPT, the first public large-scale corpus for the Brazilian Portuguese medical domain111Due storage constraints, during blind review we only release a 10K sample of the dataset, comprising 384,095 authentic question-answer pairs from real patient-doctor interactions along with metadata detailing the disease, the medical specialty, and a label identifying the question type. We provide a comprehensive analysis of the dataset, detailing its linguistic properties, such as the natural asymmetry of patient and doctor communication, and its thematic breadth, which covers over 3,200 medical topics and 104 professional specializations. To validate its practical utility, we establish strong benchmarks on a real-world medical subspecialty classification task. We demonstrate that fine-tuning a modern 1.7B parameter model on MedPT achieves a 94% F1-score on a challenging 20-class setup, confirming the dataset is a high-quality and effective resource for developing and specializing language models for the nuanced Brazilian healthcare context.
@@ 2. Related Works
However, the development and fine-tuning of such models require large, well-structured, and high-quality datasets. To address this, several datasets have been proposed to advance research in medical natural language processing. For example, Johnson et al. (2023) introduced MIMIC-IV, which focuses on describing a vast collection of de-identified electronic health records (EHR) from intensive care units, widely used for developing clinical prediction models and training NLP systems on real-world clinical documentation. Similarly, BioRED Luo et al. (2022) aimed at advancing biomedical knowledge extraction by annotating document-level relations between diverse entities like genes, diseases, and chemicals from scientific papers abstracts.
Despite these contributions, most existing datasets consist primarily of clinical or biomedical records and lack open-ended, real-world patient questions paired with verified medical answers. In response to this limitation, He et al. (2019) proposed cMedQA, which includes 54,000 pairs of physician-patient questions and answers in Chinese, sourced from online health forums, focusing on non-restricted, open-ended questions. Following this line of research, MedRedQA Nguyen et al. (2023) includes 51,000 pairs of open-ended English consumer questions and verified expert answers, collected from online medical consultation forums.
In response to this data scarcity, pioneering efforts have begun to establish benchmark medical datasets for several under-resourced languages, employing a variety of curation strategies. One common approach involves collecting data from trusted online health portals. For instance, ViHealthQA Nguyen et al. (2022) contains over 10,000 medical question-answering pairs for Vietnamese sourced from credible health websites, while MAQA Abdelhay et al. (2023) represents the largest Arabic healthcare Q&A dataset, with over 430,000 questions from regional health platform. Other initiatives have utilized different sources; MedTurkQuAD İncidelen and Aydoğan (2024), a Turkish medical QA dataset, was derived from Turkish Wikipedia and national medical theses. The effort for Persian by Ghassabi et al. (2025) was even more comprehensive, introducing the MF3QA dataset and leveraging it to create Gaokerena-V1, the first open-source Persian medical language model. Together, these works illustrate a growing cross-lingual interest in expanding medical NLP resources beyond English, particularly for morphologically rich or culturally distinct languages.
@@
A recent initiative targeting this conversational gap is IaraMed Färber et al. (2025), a women’s healthcare chatbot developed using data scraped from online health platform. While a promising application of real-world medical Q&A, this work is narrowly focused on women’s health and does not cover the full spectrum of patient-doctor inquiries, and the underlying dataset was not publicly released. Thus, a critical need persists for a broad-domain, publicly accessible, native corpus built from authentic patient-doctor interactions to advance the field.
@@ 3. MedPT
This section details the multi-stage process applied in the development of MedPT dataset, designed to transform raw, user-generated content into a robust language resource. We first describe the data acquisition from an authentic patient-doctor interaction platform, followed by our extensive data cleaning and curation pipeline, which includes both general preprocessing and nuanced, column-specific refinements. We then detail an LLM-driven annotation step, where we employed a large language model to classify each query and extract a new, structured feature to enhance the dataset’s utility. Finally, to ensure fair comparison and compatibility with modern models, the entire corpus was processed using the sub-word tokenizer employed by state-of-the-art LLMs Achiam et al. (2023).
@@ 3.1. Data Acquisition
To build the MedPT dataset, we collected real medical questions and answers in Portuguese from Doctoralia222https://www.doctoralia.com.br, a widely used online platform where Brazilian patients ask health-related questions and receive responses from verified doctors. Doctoralia was chosen as data source due both its large volume of questions across a wide range of medical professions (specialties) and its verification process to ensure that only certified health professionals can respond to users’ inquiries. Additionally, all questions are completely anonymous, which guarantees the privacy and confidentiality of users’ personal information. Furthermore, the data collected from this platform are natural and user-generated, reflecting real-life inquiries from Brazilian users. This aspect makes the dataset particularly valuable for developing healthcare solutions tailored to the Brazilian context, as it captures the country’s linguistic and cultural nuances, healthcare system characteristics, and the prevalence of region-specific diseases.
To collect the questions, answers, and related data, we developed a custom web-scraping pipeline to extract information directly from the website. Each sample includes the medical topic of the question (e.g., hypertension, pregnancy, diabetes), the patient’s textual query, the doctor’s textual response, and metadata about the physician, such as profession and rating. During this collection phase, we gathered approximately 2 million raw samples.
@@ 3.2.1. Column-Specific Curation
To ensure the quality and suitability of the corpus for question-answering tasks, we implemented a multi-stage, column-specific curation protocol through all features of our dataset (Question, Answer Question Topic and Profession). This protocol combined quantitative analysis with qualitative manual inspection. We first analyzed token distributions such as mean, standard deviation (SD) and range, to identify statistical outliers, which were then manually reviewed to confirm their nature. Based on this hybrid analysis, we applied targeted procedures, including the removal of malformed or uninformative entries and the contextual enrichment of incomplete questions, as detailed below.
@@
The Question column exhibited high linguistic diversity (mean 42.5, SD 31.0), with token counts ranging from 3 to 801. Our qualitative inspection confirmed that the longest entries were valid, complex patient narratives. Conversely, the shortest entries (3–9 tokens) were typically found to be context-dependent (e.g., "What is the treatment?"), lacking sufficient information to be standalone queries. To address this ambiguity without data loss, we implemented a contextual enrichment procedure: the corresponding Question Topic was prepended to each of these 7,178 incomplete queries. This process rendered each question self-contained and preserved a significant portion of the dataset that would have otherwise been discarded.
@@
Question Topic.
For the Question Topic column, we implemented a token-count-based filtering protocol to ensure data quality. An initial analysis programmatically removed all zero-token samples. We then conducted a qualitative review of the distribution’s extremes. This inspection revealed that single-token entries (e.g., "pain") were valid medical topics and were therefore retained. Conversely, entries with the highest token counts were found to be consistently malformed or corrupted data. Based on this manual review, we established a conservative upper threshold of 23 tokens to precisely exclude these erroneous entries while preserving all valid topics. This procedure refined the column’s consistency by removing 398 samples.
Profession.
The Profession column presented a distinct curation challenge. Statistical analysis showed moderate variance (mean 5.7 tokens, SD 3.4), which our qualitative review attributed to the platform’s feature allowing professionals to list multiple specialties. We made the deliberate decision to retain these multi-specialty entries to avoid information loss, as they represent valid, rich data. However, a separate filtering process was necessary to address genuine scraping artifacts. Our analysis identified and removed non-professional data, such as location names (e.g., "Fortaleza"), numerical values, and empty strings. This targeted procedure refined the column’s integrity by excluding 62 invalid samples.
@@ 3.3. Question-Type Annotation
To provide a fine-grained, semantic understanding of user intent within the corpus, we enriched the dataset with a synthetically generated feature column classifying the type of each user question. This structural annotation is critical for analyzing the dataset’s composition (e.g., quantifying the types of user needs) and enables more targeted modeling applications. Following established practices in dataset construction Alhuzali et al. (2024); Guo et al. (2018), we employed gpt-oss-120b Agarwal et al. (2025) to annotate each user query. Through a structured prompt and definitions, the model assigned each question to one of the following seven predefined classes, which represent distinct facets of health-related inquiries:
@@ 4. Dataset Characterization
To characterize the final MedPT corpus and validate its suitability for training robust language models, we performed an extensive exploratory data analysis. This section details the dataset’s key statistical and structural properties, including an analysis of the distribution across question types and medical professions, a summary of its linguistic features, and a presentation of representative samples. The analysis underscores the corpus’s large scale and multi-domain nature, confirming its value as a comprehensive resource for the Brazilian Portuguese medical NLP community.
@@ 4.1. Dataset Overview
Question Topic: The primary medical field of the inquiry (e.g., Gynecology, General Practice, Dermatology).
Subspecialty: A finer-grained topic within the main field, often a specific disease or condition (e.g., Diabetes, Acne, Hypertension).
Profession: The specialization of the professional who authored the answer (e.g., General Practitioner, Psychologist, Nutritionist).
@@ 4.3. Diversity and Coverage
Topic Questions Answers Frequent - Questions Syphilis 2,336 2,603 Inguinal Hernia 1,914 2,826 Appendicitis 1,773 2,250 Frequent - Answers Depression 1,140 11,715 Anxiety 1,025 10,734 Bipolar Disorder 928 3,906 Frequent - Both HIV and AIDS 5,664 6,162 COVID-19 4,674 5,339
Beyond its scale, MedPT exhibits substantial thematic diversity, covering over 3,200 unique medical inquiry topics. An analysis of the topic distribution reveals wide coverage across medical domains and highlights a key characteristic of the dataset. As shown in Table 1, the most frequent topics among patient questions include infectious and clinical diseases such as HIV/AIDS, COVID-19, and Syphilis, while the topics with the highest volume of professional answers are concentrated in mental health fields. This distinction underscores the dataset’s unique value, capturing a broad spectrum of both urgent public health concerns from patients and areas of deep, sustained engagement from healthcare professionals.
Furthermore, Table 2 demonstrates that topics with few questions often received a huge amount of answers, suggesting that even rare or highly specific issues tend to stimulate extensive professional interaction. This pattern indicates that, although these topics represent a smaller portion of the dataset, they attract multiple perspectives and deeper discussions around the same underlying question.
Topic Questions Answers Food desensitization 1 39 Nightmares 1 38 Changes After Death 2 75 Relationship conflicts 3 112 Mutism 1 35 Existential distress 2 66 Difficulty in decision-making 3 96 Deep and prolonged sadness 3 95 Shyness 10 311
To facilitate a finer-grained analysis of user intent, each question was enriched with a Question Type label. The distribution of these categories, shown in Figure 1, reveals that the dataset is dominated by action-oriented inquiries. Questions concerning Treatment and Diagnosis collectively account for nearly 70% of the data, indicating a primary focus on actionable medical guidance. Concurrently, the presence of informational categories such as Epidemiology and Healthy Lifestyle confirms the dataset’s thematic breadth. This diversity is also reflected in the professional specializations of the answer providers. The dataset includes answers from professionals across 104 distinct titles, which form over 900 unique specialization combinations due to multi-specialty declarations.

### v2

@@ Abstract
While large language models (LLMs) show transformative potential in healthcare, their development remains focused on high-resource languages. This creates a critical barrier for other languages, as simple translation fails to capture unique clinical and cultural nuances, such as endemic diseases. To address this, we introduce MedPT, the first large-scale, real-world corpus of patient-doctor interactions for the Brazilian Portuguese medical domain. Comprising 384,095 authentic question-answer pairs and covering over 3,200 distinct health-related conditions, the dataset was refined through a rigorous multi-stage curation protocol that employed a hybrid quantitative-qualitative analysis to filter noise and contextually enrich thousands of ambiguous queries, resulting in a corpus of approximately 57 million tokens. We further utilize of LLM-driven annotation to classify queries into seven semantic types to capture user intent. To validate MedPT’s utility, we benchmark it in a medical specialty classification task: fine-tuning a 1.7B parameter model achieves an outstanding 94% F1-score on a 20-class setup. Furthermore, our qualitative error analysis shows misclassifications are not random but reflect genuine clinical ambiguities (e.g., between comorbid conditions), proving the dataset’s deep semantic richness. We publicly release MedPT on Hugging Face to support the development of more equitable, accurate, and culturally-aware medical technologies for the Portuguese-speaking world.
@@
MedPT: A Massive Medical Question Answering Dataset for Brazilian-Portuguese Speakers††thanks: Accepted at LREC 2026.
@@ 1. Introduction
Large language models (LLMs) are rapidly expanding their reach across diverse healthcare applications, from online patient support and clinical text classification to robust medical question answering Nazi and Peng (2024); Liu et al. (2025). However, the performance, safety, and clinical reliability of these systems are fundamentally contingent on the scale and quality of their training data Ryan Conmy et al. (2023). Consequently, datasets that lack demographic representation, contain inherent biases, or are insufficiently curated to handle complex medical nuances can lead to models that produce factually incorrect or systematically skewed outputs, posing significant risks and undermining clinical trust Khan et al. (2021); Schwabe et al. (2024). Compounding this challenge, current progress and resource development are overwhelmingly concentrated in high-resource languages, particularly English, a linguistic bias that creates a critical barrier for other major world languages, severely limiting the development, validation, and equitable deployment of these promising technologies to large global populations Enshaei and Naderkhani (2024).
This linguistic and data-centric gap is clearly reflected in the literature, where significant research investment in medical text curation has centered on English Jin et al. (2019); Pal et al. (2022); Singhal et al. (2022); Preston et al. (2023), Chinese He et al. (2019); Zhang et al. (2018); Zhu et al. (2023), or both He et al. (2020). Despite valuable initiatives have begun to extend this scope to other languages such as Arabic, Vietnamese, Persian, and Turkish Abdelhay et al. (2023); Nguyen et al. (2022); Ghassabi et al. (2025); İncidelen and Aydoğan (2024), the resource disparity remains particularly acute for Brazilian Portuguese in specialized domains, such as medical question answering.
@@
In this work, we introduce MedPT111https://huggingface.co/datasets/AKCIT/MedPT, the first public large-scale corpus for the Brazilian Portuguese medical domain, comprising 384,095 authentic question-answer pairs from real patient-doctor interactions along with metadata detailing the disease, the medical specialty, and a label identifying the question type. We provide a comprehensive analysis of the dataset, detailing its linguistic properties, such as the natural asymmetry of patient and doctor communication, and its thematic breadth, which covers over 3,200 medical conditions across 104 medical specialties. To validate its practical utility, we establish strong benchmark on a real-world medical speciality classification task, achieving 94% F1-score on a challenging 20-class setup by fine-tuning a modern 1.7B parameter model on MedPT, confirming the dataset is a high-quality and effective resource for developing and specializing language models for the nuanced Brazilian healthcare context.
@@ 2. Related Works
However, the development and fine-tuning of such models require large, well-structured, and high-quality datasets. To address this, several datasets have been proposed to advance research in medical natural language processing. For example, Johnson et al. (2023) introduced MIMIC-IV, which focuses on describing a vast collection of de-identified electronic health records from intensive care units, widely used for developing clinical prediction models and training NLP systems on real-world clinical documentation. Similarly, BioRED Luo et al. (2022) aimed at advancing biomedical knowledge extraction by annotating document-level relations between diverse entities like genes, diseases, and chemicals from scientific papers abstracts.
Despite these contributions, most existing datasets consist primarily of clinical or biomedical records and lack open-ended, real-world patient questions paired with verified medical answers. He et al. (2019) proposed cMedQA, which includes 54,000 pairs of physician-patient questions and answers in Chinese, sourced from online health forums, focusing on non-restricted, open-ended questions. Similarly, MedRedQA Nguyen et al. (2023) includes 51,000 pairs of open-ended English consumer questions and verified expert answers, collected from online medical consultation forums.
Moreover, pioneering efforts have begun to establish benchmark medical datasets for several under-resourced languages, employing a variety of curation strategies. One common approach involves collecting data from trusted online health portals. For instance, ViHealthQA Nguyen et al. (2022) contains over 10,000 medical question-answering pairs for Vietnamese sourced from credible health websites, while MAQA Abdelhay et al. (2023) represents the largest Arabic healthcare Q&A dataset, with over 430,000 questions from regional health platform. Other initiatives have utilized different sources; MedTurkQuAD İncidelen and Aydoğan (2024), a Turkish medical QA dataset, was derived from Turkish Wikipedia and national medical theses. The effort for Persian by Ghassabi et al. (2025) was even more comprehensive, introducing the MF3QA dataset and leveraging it to create Gaokerena-V1, the first open-source Persian medical language model. Together, these works illustrate a growing cross-lingual interest in expanding medical NLP resources beyond English, particularly for morphologically rich or culturally distinct languages.
@@
A recent initiative targeting this conversational gap in Portuguese is IaraMed Färber et al. (2025), a women’s healthcare chatbot developed using data scraped from online health platform. While a promising application of real-world medical Q&A, this work is narrowly focused on women’s health and does not cover the full spectrum of patient-doctor inquiries, and the underlying dataset was not publicly released. Thus, a critical need persists for a broad-domain, publicly accessible, native corpus built from authentic patient-doctor interactions to advance the field.
@@ 3. MedPT
This section details the multi-stage process applied in the development of MedPT dataset, designed to transform raw, user-generated content into a robust language resource. We first describe the data acquisition from an authentic patient-doctor interaction platform, followed by our extensive data cleaning and curation pipeline, which includes both general preprocessing and nuanced, column-specific refinements. We then detail an LLM-driven annotation step, where we employed a large language model to classify each user query and identify its question type (e.g., treatment, diagnosis). to enhance the dataset’s utility. Finally, to ensure a fair comparison and compatibility with modern models, all metrics were extracted using the sub-word tokenizer adopted by state-of-the-art LLMs Achiam et al. (2023).
@@ 3.1. Data Acquisition
To build the MedPT dataset, we collected real medical questions and answers in Portuguese from Doctoralia222https://www.doctoralia.com.br, a widely used online platform where Brazilian patients ask health-related questions and receive responses from verified doctors. Doctoralia was chosen as data source due both its large volume of questions across a wide range of medical specialties and its verification process to ensure that only certified health professionals can respond to users’ inquiries. Additionally, all questions are completely anonymous, which guarantees the privacy and confidentiality of users’ personal information. Furthermore, the data collected from this platform are natural and user-generated, reflecting real-life inquiries from Brazilian users. This aspect makes the dataset particularly valuable for developing healthcare solutions tailored to the Brazilian context, as it captures the country’s linguistic and cultural nuances, healthcare system characteristics, and the prevalence of region-specific diseases.
To collect the questions, answers, and related data, we developed a custom web-scraping pipeline to extract information directly from the website. Each sample includes the medical condition of the question (e.g., hypertension, pregnancy, diabetes), the patient’s textual query, the doctor’s textual response, and metadata about the physician, such as specialty and rating. During this collection phase, we gathered approximately 2 million raw samples.
@@ 3.2.1. Column-Specific Curation
To ensure the quality and suitability of the corpus for question-answering tasks, we implemented a multi-stage, column-specific curation protocol through all features of our dataset (Question, Answer Condition and Medical Specialty). This protocol combined quantitative analysis with qualitative manual inspection. We first analyzed token distributions such as mean, standard deviation (SD) and range to identify statistical outliers, which were then manually reviewed to confirm their nature. Based on this hybrid analysis, we applied targeted procedures, including the removal of malformed or uninformative entries and the contextual enrichment of incomplete questions, as detailed below.
@@
The Question column exhibited high linguistic diversity (mean 42.5, SD 31.0), with token counts ranging from 3 to 801. Our qualitative inspection confirmed that the longest entries were valid, complex patient narratives. Conversely, the shortest entries (3–9 tokens) were typically found to be context-dependent (e.g., "What is the treatment?"), lacking sufficient information to be standalone queries. To address this ambiguity without data loss, we implemented a contextual enrichment procedure: the corresponding Condition was prepended to each of these 7,178 incomplete queries. This process rendered each question self-contained and preserved a significant portion of the dataset that would have otherwise been discarded.
@@
Condition.
For the Condition column, we implemented a token-count-based filtering protocol to ensure data quality. An initial analysis programmatically removed all zero-token samples. We then conducted a qualitative review of the distribution’s extremes. This inspection revealed that queries containing very few tokens (e.g., "anxiety") were valid medical conditions and were therefore retained. Conversely, entries with the highest token counts were found to be consistently malformed or corrupted data. Based on this manual review, we established a conservative upper threshold of 23 tokens to precisely exclude these erroneous entries while preserving all valid conditions. This procedure refined the column’s consistency by removing 398 samples.
Medical Specialty.
The Medical Specialty column presented a distinct curation challenge. Statistical analysis showed moderate variance (mean 5.7 tokens, SD 3.4), which our qualitative review attributed to the platform’s feature allowing professionals to list multiple specialties. We made the deliberate decision to retain these multi-specialty entries to avoid information loss, as they represent valid, rich data. However, a separate filtering process was necessary to address genuine scraping artifacts. Our analysis identified and removed non-specialty data, such as location names (e.g., "Fortaleza"), numerical values, and empty strings. This targeted procedure refined the column’s integrity by excluding 62 invalid samples.
@@ 3.3. Question-Type Annotation
To provide a fine-grained, semantic understanding of user intent within the corpus, we enriched the dataset with a synthetically generated feature column classifying the type of each user question. This structural annotation is critical for analyzing the dataset’s composition (e.g., quantifying the types of user needs) and enables more targeted modeling applications. Following established practices in dataset construction Alhuzali et al. (2024); Guo et al. (2018), we employed gpt-oss-120b Agarwal et al. (2025) to annotate each user query. Through a structured prompt, avalable in Appendix A, the model assigned each question to one of the following seven predefined classes, which represent distinct facets of health-related inquiries:
@@ 4. Dataset Characterization
To characterize the final MedPT corpus and validate its suitability for training robust language models, we performed an extensive exploratory data analysis. This section details the dataset’s key statistical and structural properties, including an analysis of the distribution across question types and medical specialties, a summary of its linguistic features, and a presentation of representative samples. The analysis underscores the corpus’s large scale and multi-domain nature, confirming its value as a comprehensive resource for the Brazilian Portuguese medical NLP community.
@@ 4.1. Dataset Overview
Condition: A finer-grained topic within the main field, often a specific disease or condition (e.g., Diabetes, Acne, Hypertension).
Medical Specialty: The specialization of the professional who authored the answer (e.g., General Practitioner, Psychologist, Nutritionist).
@@ 4.2. Quantitative Analysis
Condition Questions Answers Frequent - Questions Syphilis 2,336 2,603 Inguinal Hernia 1,914 2,826 Appendicitis 1,773 2,250 Frequent - Answers Depression 1,140 11,715 Anxiety 1,025 10,734 Bipolar Disorder 928 3,906 Frequent - Both HIV and AIDS 5,664 6,162 COVID-19 4,674 5,339
@@ 4.3. Diversity and Coverage
Beyond its scale, MedPT exhibits substantial thematic diversity, covering over 3,200 unique medical conditions. An analysis of the condition distribution reveals wide coverage across medical domains and highlights a key characteristic of the dataset. As shown in Table 1, the most frequent conditions among patient questions include infectious and clinical diseases such as HIV/AIDS, COVID-19, and Syphilis, while the conditions with the highest volume of professional answers are concentrated in mental health fields. This distinction underscores the dataset’s unique value, capturing a broad spectrum of both urgent public health concerns from patients and areas of deep, sustained engagement from healthcare professionals.
Furthermore, Table 2 demonstrates that conditions with few questions often received a huge amount of answers, suggesting that even rare or highly specific issues tend to stimulate extensive professional interaction. This pattern indicates that, although these conditions represent a smaller portion of the dataset, they attract multiple perspectives and deeper discussions around the same underlying question.
Condition Questions Answers Food desensitization 1 39 Nightmares 1 38 Changes After Death 2 75 Relationship conflicts 3 112 Mutism 1 35 Existential distress 2 66 Difficulty in decision-making 3 96 Deep and prolonged sadness 3 95 Shyness 10 311
To facilitate a finer-grained analysis of user intent, each question was enriched with a Question Type label. The distribution of these categories, shown in Figure 1, reveals that the dataset is dominated by action-oriented inquiries. Questions concerning Treatment and Diagnosis collectively account for nearly 70% of the data, indicating a primary focus on actionable medical guidance. Concurrently, the presence of informational categories such as Epidemiology and Healthy Lifestyle confirms the dataset’s thematic breadth. This diversity is also reflected in the specialties of the answer providers, including answers from professionals across 104 distinct specialties, which form over 900 unique specialization combinations due to multi-specialty declarations.